# Weakly-supervised diagnosis identification from Italian discharge letters

### v1

@@ Abstract
Objective: Identifying the diagnosis that determines a patient’s hospitalization or Emergency Room admission is crucial for cohort selection and epidemiological analyses. Discharge letters are the only type of data consistently produced across hospitals. Recognizing diseases from discharge letters is a classic document classification problem, typically requiring supervised learning. However, manual annotation of large datasets of discharge letters is uncommon since it is extremely time-consuming. We propose a novel weakly-supervised pipeline to recognize diseases from Italian discharge letters.
Methods: Our Natural Language Processing pipeline is based on a fine-tuned version of the Italian Umberto model. The pipeline extracts diagnosis-related sentences from a subset of letters and applies a two-level clustering using the embeddings generated by the fine-tuned Umberto model. These clusters are summarized and those mapped to the diseases of interest are selected as weak labels. Finally, the same BERT-based model is trained using these weak labels to detect the targeted diseases.
Results: Results on a case study related to the identification of bronchiolitis in a dataset of 33’176 Italian discharge letters from multiple hospitals in the Veneto Region demonstrate the potential of our method, with an AUC of 77.7 % and an F1-Score of 75.1 % on manually annotated labels, improving compared to other non-supervised methods and with a limited loss compared to fully supervised methods. Results are robust to the cluster selection and the identified clusters highlight the potential to recognize a variety of diseases.
Conclusions: This study demonstrates the feasibility of successfully addressing the challenge of diagnosis identification from Italian discharge letters in the absence of labelled data. Our weakly-supervised pipeline showed strong performance and robustness, and its flexibility allows for easy adaptation to various diseases. This approach offers a scalable solution for clinical text classification, significantly reducing the need for manual annotation while maintaining good accuracy.
@@ 1 Introduction
Electronic Health Records contain a vast and continuously growing amount of patient information, but a significant portion of this data remains unstructured in formats like clinical notes, reports, discharge letters and referrals. Despite containing valuable information, this unstructured data is often underutilized in the downstream analysis due to their complexity [1, 2]. While Natural Language Processing (NLP) has made significant advancements in extracting and classifying information from text [3, 4], applying these techniques to healthcare data presents additional challenges, such as a specialized lexicon, a high prevalence of ambiguous abbreviations, frequent misspellings, and a lack of standard templates across hospitals [5, 6]. Additionally, most NLP tools are developed for high-resource languages, making them less effective for languages like Italian, where linguistic resources are scarce [7].
A critical piece of information in medical documents is the reason for a patient’s hospital admission or discharge, typically a disease diagnosis or a description of the patient’s symptoms. This information is critical for both cohort selection and epidemiological analysis. Although an international code system for diseases and symptoms, the ICD-CM [8], exists, physicians often prefer to provide a textual description of their diagnoses. Textual descriptions are faster to write, more detailed, and are often used because many medical document templates do not include structured fields for disease codes [9, 10]. Even when present, it has been observed that ICD codes are not completely reliable [11, 12, 13]. Consequently, identifying patients admitted for specific diseases or symptoms often requires manual inspection of these texts, which is labour-intensive and impractical at scale.
In Italy, every hospitalization or Emergency Room (ER) admission generates a discharge letter. Depending on the type of examinations and the hospital, more structured data may also be recorded. A discharge letter should always contain information about the diseases that affect the patient, but the complexity of these documents and the lack of standard templates pose a significant challenge for automatic data extraction. To address this issue, this study aims to develop an NLP pipeline capable of extracting structured information from the unstructured text of Italian discharge letters.
Traditional approaches to this problem typically frame it as a supervised classification task, which requires a substantial amount of annotated data to perform well [14]. However, annotating thousands of discharge letters is time-consuming, especially for clinicians who are already overworked. This highlights a significant gap in the field: the need for methods that can effectively utilize large, unannotated datasets without the burden of manual labelling.
While there have been a few studies on NLP applied to Italian healthcare documents, none have focused specifically on discharge letters. Lanera et al. [15, 16] worked on supervised models to identify cases of varicella, using textual data from specific children’s health project dataset. Similarly, Sciannameo et al. [17] used a BERT-based model to predict rehospitalizations for elderly people from the text of medication prescriptions and hospitalization diagnoses. Hammami et al.  [18] classified pathology reports for cancer morphology, developing a rule-based system. Other studies have concentrated on extracting events and entities, such as Viani et al. [19, 20], who analyzed cardiology reports using ontologies and recurrent neural networks.
Significant research on discharge letters has been done in the English language, often using the MIMIC dataset [21, 22] (see [23, 24, 25, 26] for some examples).
Research has also been conducted on discharge letters in other under-resourced languages, such as Bulgarian [27, 28], Dutch [29] and Swedish [30].
All these studies have relied either on large annotated datasets or on the development of complex, problem-specific rule-based systems. In contrast, our approach eliminates the dependency on manually annotated data and offers a more generalizable solution than traditional rule-based systems.
The main contribution of this paper is the first NLP-based pipeline for the weakly-supervised classification of Italian discharge letters, offering a scalable solution to a long-standing challenge in healthcare data management.
The results presented in this paper pertain to a clinically relevant case study aimed at identifying bronchiolitis, one of the most relevant diseases that can determine hospitalizations in young children [31]. An automated model for this purpose can significantly aid in patient cohort selection for research, reducing the manual workload for physicians. The dataset for this case study comprises 33’176 discharge letters related to children from 44 different hospitals in the Veneto region, one of the most populated Italian regions.
The rest of the paper is structured as follows: in Section 2, we discuss the methodology we developed and the data of our case study; in Section 3, we present the results; in Section 4 we discuss them, together with possible future developments, and in Section 5 we summarize the main findings.
Statement of significance
Problem Identifying diagnoses from discharge letters is relevant for both cohort selection and epidemiological analysis but requires large annotated datasets, rarely available in this domain.
What is Already Known Supervised learning methods for diagnosis identification are well-studied, with state-of-the-art methods based on transformers. Unsupervised methods are mainly rule-based algorithms that lack generalizability.
What this Paper Adds This paper introduces a weakly-supervised NLP pipeline that eliminates the need for manual annotations, offering a scalable solution for diagnosis identification. The approach is validated on a case study with a large dataset of Italian discharge letters and demonstrates strong performance and potential generalizability.
@@ 2 Methods
In this section, we describe in detail the novel pipeline we propose for the weakly-supervised diagnosis identification. The pipeline is sketched in Figure 1 and it can be divided into three main steps:
@@
    BPT - Pre-training of a BERT-based model for Italian medical documents.
@@
Step 0 (BPT) is described in Section 2.0.1. It aims to build a model able to produce embeddings of Italian medical documents that are exploited by both Step 1 (SAL) and Step 2 (WLC), described in Section 2.0.2 and 2.0.3.
Figure 1: Pipeline schema. Diagram of the full pipeline for weakly-supervised classification of discharge letters
2.0.1 Pre-training of a BERT-based model for Italian medical documents (BPT)
@@
Statistical and machine learning models for textual data require a mathematical representation of the text. There are three main types of possible representations: vocabulary-based, static neural network-based embeddings and contextual neural network-based embeddings. Vocabulary-based representations produce a high dimensional vector whose size is equal to the corpus vocabulary size and whose content is sparse. Their main issue is that they cannot capture the semantic relationships between words, i.e., the fact that different words may have the same or similar meaning, possibly also depending on the context in which they appear. These issues are overcome by neural network-based embeddings. In particular, static word embeddings, such as Word2Vec [32], produce embedding vectors that are permanently fixed for each term. In contrast, contextual embeddings are related to the context in which the word appears, capturing more semantics at the price of requiring an inference through a neural network each time the embedding has to be computed. The state-of-the-art model for contextual embeddings is BERT (Bidirectional Encoder Representations from Transformers) [33], a deep learning model based on the transformer architecture [34]. Different versions of BERT have been developed for different languages and domains. In particular, for the English language, there are various versions of BERT related to the medical domain, such as BioBERT [35] and ClinicalBERT [36].
For the Italian language, the only available versions of BERT are trained on general-domain documents crawled from the web [37]. Thus, we fine-tuned the last Italian version of BERT, Umberto, using the E3C Corpus [38], a multi-lingual dataset composed of documents in the medical domain, particularly clinical narratives from various sources: journal papers, admission tests for specialities in medicine, patient information leaflets for medicines, and abstracts of theses in medical science. The dataset contains documents in five languages; a summary of its composition is provided in Supplementary Material A3. This is the only publicly available dataset that includes Italian medical documents. To utilize the entire dataset, which amounts to 564 MB, we automatically translated the non-Italian documents using Google Translate’s APIs (Step 0.1 in Figure 1). We further pre-trained the Umberto model with this dataset, adopting the usual masked-word prediction task (Step 0.2 in Figure 1). This additional pre-training of the Umberto model over the E3C corpus involved 3 epochs of training (until the validation perplexity did not improve further) with a learning rate of $2\cdot 10^{-5}$ and weight decay of $0.01$, using 20% of the data as a validation set for early stopping.
2.0.2 Semi-automatic labelling pipeline (SAL)
@@
We aim to develop a classification system for discharge letters that can work on an unlabelled dataset. To do so, we aim to construct weak labels that a supervised classification model can subsequently use. This section presents a methodology to construct weak labels by extracting diagnosis strings from discharge letters and clustering them. This allows the creation of groups of diagnoses related to the same disease or symptoms that can be exploited as labels. This step corresponds to Step 1 (SAL) of Figure 1.
@@
Our approach combines regular expressions for extracting diagnosis strings and a customized clustering pipeline based on the BERT model.
In the following paragraphs, we describe in detail these sub-steps.
Extraction of diagnosis strings
@@
The first step, corresponding to block 1.1 in Figure 1, is the extraction of diagnosis strings. Italian discharge letters do not follow a standardized template, as each hospital has its own, and even different departments within the same hospital may personalize the template further. Nevertheless, we have observed that these templates often include a designated section for the diagnosis, where the doctor should report the patient’s admission diagnosis. This diagnosis indication should contain, in most cases, the necessary information to determine the label for the discharge summary. However, there are three main challenges in handling this diagnosis indication:
@@
The first point is the primary reason why our pipeline focuses on classifying the text of a discharge letter rather than just the diagnosis string. To address the second point, we employ regular expressions to detect the portion of text where they are present. Regular expressions provide a powerful and flexible approach for pattern matching, making them suitable for capturing a template field with limited variations, such as the diagnosis indication in our dataset. We reported in Supplementary Material C the expressions we defined to extract the diagnosis strings.
@@
Diagnosis Clustering
Once the diagnosis strings are extracted, we use a tailored clustering pipeline to group them meaningfully, aiming at identifying clusters related to the various types of diseases and symptoms. This step corresponds to block 1.2 in Figure 1 and is expanded in Figure 2.
Figure 2: Clustering section of the pipeline. Detailed diagram of the clustering pipeline for diagnosis strings, corresponding to Step 1.2 of the full pipeline depicted in Figure 1.
Punctuation removal, digits removal and abbreviation expansion are applied to the diagnosis strings (step 1.2.1). This pre-processing aims to facilitate the subsequent construction of a text representation that can be representative of the semantics of the text. In particular, punctuation signs and digits are irrelevant for discriminating between diagnoses and can be safely removed. The problem of abbreviations is a common problem in medical documents: a wide variety of acronyms exist in the medical domain, and their meaning can also be context-dependent. We have extracted all the words with two and three letters from our dataset of diagnosis strings and expanded those that we found to be medical abbreviations.
After this pre-processing, the pipeline needs to construct an embedding of the text that a clustering algorithm will subsequently use. This step (1.2.2) exploits the BERT model we discussed in Section 2.0.1 (BPT).
The BERT-based models produce embedding vectors whose size is 768, making it necessary to apply a dimensionality reduction technique over these embeddings to use them as input for a clustering algorithm (Step 1.2.3).
After this step, a clustering algorithm can be applied to the resulting embedding vectors (Step 1.2.4). We selected HDBSCAN since it provides many relevant advantages: it does not require defining a priori the number of clusters, it does not necessarily cluster all the data, and it can identify clusters of varying shapes and sizes.
To effectively group strings related to different diseases and symptoms, the clustering must prioritize the intra-cluster similarity against the inter-cluster dissimilarity. This has the drawback of creating multiple clusters associated with each disease/symptom that are very compact internally, but that should be merged. A second-level clustering can help merge these clusters that are initially separate, simplifying the identification of the weak labels. To this purpose, a suitable representation for clusters needs to be produced. We employ a keyword representation in which, for each cluster, we extract words that appear with significant frequency in the cluster and do not appear with considerable frequency over the entire dataset (Step 1.2.5). These keywords can be represented with BERT embeddings in the same way as the original diagnosis strings and used for subsequent clustering (Step 1.2.6).
Mapping clusters to diseases

### v2

# Automatic identification of diagnosis from hospital discharge letters via weakly-supervised Natural Language Processing

@@ Abstract
Identifying patient diagnoses from hospital discharge letters is essential to enable large-scale cohort selection and epidemiological research, but traditional supervised learning approaches rely on extensive manual annotation, which is often impractical for large textual datasets. In this study, we present a novel weakly supervised Natural Language Processing pipeline designed to classify Italian discharge letters without requiring manual labelling. After extracting diagnosis-related sentences, the method leverages a transformer-based model with an additional pre-training on Italian medical documents to generate semantic embeddings. A two-level clustering procedure is then applied to these embeddings, and the resulting clusters are mapped to the diseases of interest to derive weak labels for a subset of data, which are subsequently used to train a transformer-based classifier. We evaluate the approach on a real-world case study focused on detecting bronchiolitis in a corpus of 33,176 Italian discharge letters of children admitted to 44 emergency rooms or hospitals in the Veneto Region between 2017 and 2020. The pipeline achieves an area under the curve (AUC) of 77.68% ($\pm 4.30\%)$ and an F1-score of 78.14% ($\pm 4.89\%$) against manually annotated data. Its performance surpasses other unsupervised methods and approaches that of fully supervised models, while maintaining robustness to cluster selection and a promising generalizability across different disease types. It allows saving approximately 3 minutes of expert time per discharge letter, corresponding to more than 1,500 hours for a dataset of the size used in our case study. This study demonstrates the feasibility of a weakly-supervised strategy for identifying diagnoses from Italian discharge letters. The pipeline achieves strong performance, is adaptable to various diseases, and offers a scalable solution for clinical text classification, reducing the need for manual annotation while maintaining reliable accuracy.
@@ 1 Introduction
Electronic Health Records contain a vast and continuously growing amount of patient information, but a significant portion of this data remains unstructured in formats like clinical notes, reports, discharge letters and referrals. Despite containing valuable information, this unstructured data is often underutilized in the downstream analysis due to their complexity and heterogeneity [tayefi2021challenges, cowie2017electronic]. While Natural Language Processing (NLP) has made significant advancements in extracting and classifying information from text [nadkarni2011natural, khurana2023natural], applying these techniques to clinical narratives remain challenging. Medical documents are characterized by a highly specialized lexicon, with ambiguous abbreviations and frequent misspellings, a lack of standard templates across hospitals and a limited amount of manually labelled data [hossain2023natural, friedman2014natural, iroju2015systematic]. Additionally, most NLP tools are optimized for high-resource languages such as English, limiting their transferability to languages like Italian, where datasets and domain-specific linguistic resources are scarce [neveol2018clinical].
Among the various types of clinical documents, hospital discharge letters play a particularly important role: they summarize the patient’s condition, diagnoses, and treatments, representing a relevant source of diagnostic information for both clinical follow-up and population-based studies. Although international standard such as the ICD-CM [world2016international] exists for encoding diagnoses, physicians often prefer to provide textual descriptions of their diagnoses because they are faster to write, more expressive, and often the only option available in document templates [walsh2004clinician, millares2021consultation]. Even when ICD-CM codes are included, previous research has shown that they can be incomplete or inaccurate [walsh2018icd, cozzolino2018accuracy, cipparone2015inaccuracy]. Automatic methods capable of identifying diagnoses directly from unstructured discharge letters are therefore crucial for cohort selection and epidemiological research [dong2022automated], reducing the need for manual review, which is labor-intensive and impractical at scale.
In Italy, every hospital or emergency room (ER) admission generates a discharge letter that should contain at least one diagnosis describing the reason for admission or discharge. However, these letters vary widely in format and terminology, often mixing narrative and structured components. Developing robust NLP systems for such data is challenging: traditional approaches rely on supervised classification models that require a substantial amount of manually annotated data [mehrafarin2022importance], an effort that is costly and time-consuming for clinicians. This highlights a significant gap in the field: the need for methods that can effectively exploit large, unannotated datasets to extract clinically meaningful information without the burden of manual labelling.
Previous studies applying NLP to Italian clinical texts have not addressed discharge letters. Lanera et al. [lanera2020use, lanera2022deep] worked on supervised models to identify varicella cases from notes of family pediatricians, while Sciannameo et al. [veronica2022deep] used a BERT-based model to predict rehospitalizations for elderly patients from medication prescriptions and hospitalization records. Hammami et al. [hammami2021automated] classified pathology reports for cancer morphology using a rule-based approach, and Viani et al. [viani2018information, viani2019supervised] extracted entities from cardiology reports through ontologies and recurrent neural networks. Most international work on discharge letters has been conducted in English using resources such as the MIMIC database [johnson2016mimic, johnson2023mimic] (see [edin2023automated, li2020icd, li2018automated, falter2024using] for some examples), but research has also been conducted in other European languages with limited resources, such as Bulgarian [boytcheva2011automatic, velichkov2020automatic], Dutch [bagheri2020automatic, van2024using], Czech [lenc2024czech] and Swedish [remmer2021multi]. These studies, however, all relied on large annotated datasets, making them difficult to reproduce in settings where manual labelling is infeasible.
Recent work has explored weakly-supervised approaches to mitigate the need for large annotated clinical corpora. Most of these methods rely on task-specific heuristics, such as disease-related keywords, regular expressions, or the transfer of labels from structured data like diagnostic codes or medication records [sanyal2022weakly, wang2019clinical, cusick2021using, wang2022leveraging, humbert2022strategies]. While effective within narrowly defined clinical settings, such strategies are difficult to extend to new diseases, domains, or institutions, as they depend on predefined lexical patterns or on the availability of reliable structured fields.
To overcome these limitations, we propose a weakly-supervised NLP pipeline for automatically identifying patient diagnoses from Italian hospital discharge letters. Our approach eliminates the dependency on manually annotated data and offers a more generalizable solution than previous rule-based methods.
The pipeline first identifies diagnosis-related sentences from a subset of documents and represents them through contextual embeddings generated by a transformer model, previously pre-trained with public data in Italian related to the medical domain. These embeddings are then grouped through a two-level clustering procedure, after which the resulting clusters are semantically summarized and mapped to diseases of interest to produce weak labels. The weakly labelled data are finally used to train a transformer-based classifier capable of recognizing those diseases directly from discharge letters. This design enables large-scale, disease-agnostic classification of clinical narratives while avoiding the costs of manual labelling.
We evaluate the proposed framework through a real-world case study focused on bronchiolitis, one of the leading causes of hospitalizations in young children [zorc2010bronchiolitis]. The dataset includes 33,176 discharge letters of pediatric patients admitted to 44 emergency rooms or hospitals across the Veneto Region (Italy) between 2017 and 2020, collected through the Pedianet network [cantarutti2021pedianet]. To the best of our knowledge, there are no prior studies addressing the identification of bronchiolitis directly from medical text, whereas existing works have relied on structured data for diagnosis prediction [durani2008clinical, gebremedhin2022developing, vartiainen2023risk].
In our evaluation, we test the weakly-supervised pipeline against manually labelled data, considering different classification backbones and comparing its performance with both fully unsupervised and fully supervised approaches. The proposed method achieves competitive performance, outperforming other unsupervised techniques and showing only a limited gap compared with supervised classifiers. Its performance is robust to weak labels selection, and the analysis of the intermediate steps indicates a strong potential for generalization across diseases.
@@ 2 Methods
Figure 1: a. Diagram of the full pipeline for weakly-supervised classification of discharge letters. Step 0 (TPT) consists in additional pre-training of a transformer-based model on a dataset of medical documents in Italian, including documents auomatically translated into Italian. Step 1 (SAL) is the core of the pipeline, consisting in the automatic extraction of diagnosis strings from the letters, their clustering (exploiting the model from TPT) and the mapping of the clusters to the disease labels. This procedure leads to a weakly-labelled subset of data that is used in Step 2 (WLC) to train a weakly-supervised classificaton model. b. Detailed diagram of the clustering step for diagnosis strings, corresponding to Step 1.2 of the full pipeline. The extracted diagnoses are embedded with a transformer-based model, and, after a dimensionality reduction step, clustered. Clusters are summarized with relevant keywords and the clustering is repeated on these clusters’ representations.
In this section, we describe in detail the pipeline we propose for the weakly-supervised diagnosis identification. The pipeline, depicted in Figure 1, can be divided into three main steps:
@@
    TPT - Pre-Training of a Transformer-based model for Italian medical documents.
@@
The next sections describe in details the three steps.
2.1 Pre-training of a Transformer-based model for Italian medical documents (TPT)
@@
Statistical and machine learning models for textual data require a mathematical representation of the text. There are three main types of possible representations: vocabulary-based, static neural network-based embeddings and contextual neural network-based embeddings. Vocabulary-based representations produce a high dimensional vector whose size is equal to the corpus vocabulary size and whose content is sparse. Their main issue is that they cannot capture the semantic relationships between words, i.e., the fact that different words may have the same or similar meaning, possibly also depending on the context in which they appear. These issues are overcome by neural network-based embeddings. In particular, static word embeddings, such as Word2Vec [mikolov2013efficient], produce embedding vectors that are permanently fixed for each term. In contrast, contextual embeddings are related to the context in which the word appears, capturing more semantics at the price of requiring an inference through a neural network each time the embedding has to be computed. The state-of-the-art models for generating contextual embeddings are based on the transformer architecture [vaswani2017attention]. Among them, BERT (Bidirectional Encoder Representations from Transformers) [devlin2019bert] and its subsequent variants, such as RoBERTa [liu2019roberta], have become foundational for a wide range of NLP tasks. Different versions of these models have been further specialized for specific languages and domains. In particular, for the English language, there are multiple models for the biomedical domain, such as BioBERT [lee2020biobert] and ClinicalBERT [alsentzer2019publicly].
For the Italian language, available transformer models equivalent to BERT are mostly trained on general-domain web data [tamburini2020bertology], with limited exposure to biomedical or clinical terminology. The E3C corpus [magnini2020e3c] is a multilingual resource that includes Italian documents from diverse medical sources, such as journal articles, specialty-school admission tests, patient information leaflets, and medical-science thesis abstracts, providing valuable clinical context for domain adaptation. The dataset contains documents in five languages; a summary of its composition is provided in A.3.
This is the only publicly available dataset that includes Italian medical documents. With the aim of better capturing domain-specific semantics, we used this data to perform an additional pre-training of the transformer-based models we included in our pipeline (Step 0.2 in Figure 1). To utilize the entire dataset, which amounts to 564 MB, we automatically translated the non-Italian documents using Google Translate’s APIs (Step 0.1 in Figure 1).
2.2 Semi-automatic labelling pipeline (SAL)
@@
Developing a weakly-supervised classification pipeline for discharge letters requires a procedure to derive weak-labels for at least a subset of data. Weak-labels are labels that can be automatically associated to the data but that might contain some noise. This section presents a methodology to construct these weak labels by extracting diagnosis strings from discharge letters and clustering them. This allows the creation of groups of diagnoses related to the same disease or symptoms that can be exploited as labels. This step corresponds to Step 1 (SAL) of Figure 1.
@@
Our approach combines regular expressions for extracting diagnosis strings and a customized clustering pipeline based on a transformer model. In the following sections, we describe in detail these sub-steps.
2.2.1 Extraction of diagnosis strings
@@
The first step (1.1 in Figure 1.a) is the extraction of diagnosis strings. Italian discharge letters do not follow a standardized template, as each hospital has its own, and even different departments within the same hospital may personalize the template further. Nevertheless, we have observed that these templates often include a designated section for the diagnosis, where the doctor should report the patient’s admission diagnosis. This diagnosis indication should contain, in most cases, the necessary information to determine the label for the discharge summary. However, there are three main challenges in handling this diagnosis indication:
@@
The first point is the primary reason why our pipeline focuses on classifying the text of a discharge letter rather than just the diagnosis string. To address the second point, we employ regular expressions to detect the portion of text where they are present. Regular expressions provide a powerful and flexible approach for pattern matching, making them suitable for capturing a template field with limited variations, such as the diagnosis indication in our dataset. C reports the set of regular expressions used to extract the diagnosis strings.
@@
2.2.2 Diagnosis Clustering
Once the diagnosis strings are extracted, we use a tailored clustering pipeline to group them meaningfully, aiming at identifying clusters related to the various types of diseases and symptoms. This corresponds to Step 1.2 in panel a of Figure 1 and is expanded in panel b of the same figure.
Minor pre-processing steps, including the removal of punctuation, dates, and time expressions, are applied to facilitate the subsequent construction of a text representation that captures the semantic content of the document, while excluding elements that are not informative for discriminating among diagnoses (Step 1.2.1).
After this pre-processing, the pipeline needs to construct an embedding of the text that a clustering algorithm will subsequently use. This step (1.2.2) exploits the transformer model we discussed in Section 2.1.
The transformer-based model produces embedding vectors whose size is 768, making it necessary to apply a dimensionality reduction technique (PCA) over these embeddings to use them as input for a clustering algorithm (Step 1.2.3).
After this step, a clustering algorithm can be applied to the resulting embedding vectors (Step 1.2.4). We selected HDBSCAN [campello2013density] because it does not require specifying the number of clusters a priori, allows leaving part of data unclustered, and can identify clusters with heterogeneous shapes and densities. These properties are well suited to transfromer-based text embeddings, which often lie on complex, nonlinear manifolds and may form clusters with varying densities, irregular geometries, and overlapping boundaries. Several commonly used clustering algorithms, such as K-Means, Gaussian Mixture Models, and standard hierarchical clustering, are less suitable for this type of data [asyaky2021improving, yeasmin2022transformer], as further supported by a comparative analysis reported in D.
To effectively group strings related to different diseases and symptoms, the clustering must prioritize the intra-cluster similarity against the inter-cluster dissimilarity. This has the drawback of creating multiple clusters associated with each disease/symptom that are very compact internally, but that should be merged. A second-level clustering can help merge these clusters that are initially separate, simplifying the identification of the weak labels. To this purpose, a suitable representation for clusters needs to be produced. We employ a keyword representation in which, for each cluster, we extract words that appear with significant frequency in the cluster and do not appear with considerable frequency over the entire dataset (Step 1.2.5). These keywords can be represented with transformer embeddings in the same way as the original diagnosis strings and used for subsequent clustering (Step 1.2.6).
2.2.3 Mapping clusters to diseases